OpenStudy (anonymous):

What are the solutions to the equation x² = 625? Choose all answers that are correct. A. –312.5 B. –25 C. 25 D. 312.5

OpenStudy (anonymous):

b,c or just b

OpenStudy (anonymous):

C becuz square root of 625 is 25 :)

OpenStudy (anonymous):

So just C

OpenStudy (anonymous):

yup

OpenStudy (abhisar):

Hey there! \(\tt \sqrt{x^2}=\pm x\)

OpenStudy (abhisar):

So, \(\sf \sqrt{625} = \pm 25\)